# TimeBridge: Better Diffusion Prior Design with Bridge Models for Time Series Generation

### v2

@@ 4.1. Data- and Time-dependent Priors for Time Series
To investigate RQ1, we suggest using data prior to having a data-dependent distribution to better approximate the data distribution. Sampling from $\mathcal{N}(\mathbf{0},\mathbf{I})$ lacks data-specific information, which can be enhanced by enforcing that the priors capture data scale and temporal dependencies. Recently, the usages of data-dependent prior have been investigated for audio (Popov et al., 2021; Lee et al., 2022) and image (Yu et al., 2024) domains. Therefore, we test the use of data-dependent prior for time series by setting the prior distribution as follows:
@@
In response to RQ2, we focus on capturing temporal dependencies in noise addition towards prior distribution. Recently, Biloš et al. (2023) injected stochastic processes instead of random noise to preserve the continuity of temporal data and Han et al. (2022) utilized the pre-trained model for prediction and set the mean of prior as the output for regression.
As a solution, we employ Gaussian processes (GPs) for prior selection, which is effective for modeling temporal data (Biloš et al., 2023; Ansari et al., 2024). GP is a distribution over function that defines the joint variability of an arbitrary pair of time points. For given sequence of timestamps $\Omega=\{1,\ldots,\tau\}$ and a function $\bm{f}_{GP}$ of timestamps that follows a GP with mean function $\bm{m}(\cdot)$ and the covariance function $\bm{\Sigma}(\cdot)$, we can draw ${\mathbf{x}}_{T}$ as follows:
@@ 4.2. Scale-preserving for Conditional Priors
To answer RQ3, we argue that the diffusion bridge is well-suited for preserving a given condition by setting the prior as the condition. The diffusion bridge has shown its effectiveness in conditional tasks such as coloring (Zhou et al., 2024), image-to-image translation (Li et al., 2023; Liu et al., 2023), and improving naive base predictions (Chen et al., 2023b; Lee et al., 2024).
Conditional generation based on trend is straightforward by (i) setting the pair-wise condition to prior ${\mathbf{x}}_{T}={\mathbf{y}}$ and (ii) training the translation from trend to data. As the expected mean of ${\mathbf{x}}_{t}$ is a linear combination of ${\mathbf{x}}_{0}$ and ${\mathbf{x}}_{T}={\mathbf{y}}$ in Equation (9), the model learns the correction starting from the trend samples, utilizing the same data scale. In contrast, standard diffusion models are unsuitable for translating conditions into synthetic data; instead, they generate new data based on a specific trend condition, typically requiring additional steps such as providing guidance or correcting intermediate data samples ${\mathbf{x}}_{t}$ during sampling. Thus, we can also eliminate the need for penalty functions or corrections for sampling.
@@
The direct case of fixed point constraints is imputation. With the mask ${\mathbf{m}}$ indicating the missing value as 0, we construct the condition ${\mathbf{c}}={\mathbf{y}}\odot{\mathbf{m}}$ for imputation where $\odot$ indicates the Hadamard product. However, unlike soft constraints that apply to all values in ${\mathbf{y}}\in\mathbb{R}^{\tau\times d}$, the condition for imputation ${\mathbf{c}}\in\mathbb{R}^{\tau\times d}$ has missing values, which should be interpolated to data-dependent values rather than zeros. Therefore, we build local linear models for condition-to-time series imputation. We adopt linear spline interpolation (De Boor, 1978) to generate a prior for fixed point constraints. In a feature-wise manner, let us consider $x\in\mathbb{R}$ as a single-dimensional input of ${\bm{x}}\in\mathbb{R}^{d}$. With a set $K$ of feature-wise observed times for $x$, the prior $x_{T}$ with the interpolation of each feature with timestamp $k$ is:
@@ 5. Related Works
Diffusion models show better performance in synthesis tasks, compared to time series generation models with VAE (Naiman et al., 2024), GANs (Yoon et al., 2019). Rasul et al. (2021) initially used diffusion networks for time series based on recurrent networks. For imputation, SSSD (Alcaraz and Strodthoff, 2023) and CSDI (Tashiro et al., 2021) considered time series imputation similar to image inpainting tasks. TimeDiff (Shen and Kwok, 2023), LDT (Feng et al., 2024), and TMDM (Li et al., 2024) focused on time series forecasting tasks. Coletta et al. (2024) investigated guiding the sampling with a penalty function called DiffTime. Our network design follows Diffusion-TS (Yuan and Qiao, 2024), which introduced a decomposition architecture to disentangle time series data into trend and seasonal components and applied a Fourier-based loss term. Recent studies also explore frequency-domain diffusion (Crabbé et al., 2024) and transformer-based alternatives (Chen et al., 2024).
@@
Several studies have brought Schrödinger bridges to time series tasks. Chen et al. (2023a) analyzed algorithmic convergence and applied the bridge to imputation. Hamdouche et al. (2023) combined kernel-based bridges with feed-forward and LSTM networks for deep hedging. Garg et al. (2024) relaxed the Schrödinger bridge by setting a geometric mixture for the prior distribution and applied their method to time series. All of these methods require iterative solvers, which are computationally expensive and incompatible with standard diffusion pipelines. Also, None of them provides a unified framework for both unconditional and conditional generation. To the best of our knowledge, our paper is the first to use diffusion bridge, extending beyond classical Schrödinger bridge formulations and providing flexible priors.
@@
Table 2. Quality comparison of unconditional time series synthetic data with length of 24. The best results are in bold and the second best are underlined. We report the average score (lower is better) and the rank of each method. The baseline experimental results (denoted by $\dagger$) are adopted from (Yuan and Qiao, 2024). We re-implement Diffusion-TS with their official code.
@@ 6.1. Experimental Setup
We assess the performance of TimeBridge using widely used time series datasets. We use two simulation datasets: Sines with 5 features of different frequencies and phases of sine functions and MuJoCo of multivariate advanced physics simulation with 14 features. For real-world datasets, we use ETT (Electricity Transformer Temperature) for long-term electric power with 7 features, Stocks of Google stock prices and volumes with 6 features, Energy, UCI dataset with appliances energy use in a low energy building with 28 features, and fMRI of the blood oxygen level-dependent functional MR imaging with 50 selected features.
For measures, we assess three metrics for unconditional time series generation on a normalized (0-1) scale. Context-Fréchet Inception Distance (Context-FID, C.-FID) score (Jeha et al., 2022) measures the distributional quality for mean and variance projected on embedding on trained TS2Vec (Yue et al., 2022). Correlational (Corr.) score (Liao et al., 2020) measures temporal dependency of real and synthetic data using the absolute error of cross-correlation. Discriminative (Disc.) score (Yoon et al., 2019) measures the performance of a classifier trained to distinguish original and synthetic data in a supervised manner. For imputation, we use the mean squared error (MSE) and mean absolute error (MAE).
Our experiments are conducted using PyTorch on four NVIDIA GeForce RTX 4090 GPUs. We report the mean and its interval of five experiments. For the detailed settings, refer to Appendix B.
@@
Table 5. Imputation results of MSE in the order of 1e-3 on the Mujoco dataset with a random mask of ratio {0.7, 0.8, 0.9}. The best results are in bold and the second best are underlined. The baseline experimental results are adopted from (Alcaraz and Strodthoff, 2023) and (Yuan and Qiao, 2024).
@@ 6.2. Unconditional Generation
We demonstrate the unconditional generation tasks with a length of 24 as baselines. We compare with TimeVAE (Desai et al., 2021), TimeGAN (Yoon et al., 2019), Cot-GAN (Xu et al., 2020), Diffwave (Kong et al., 2021), DiffTime (Coletta et al., 2024) with SSSD (Alcaraz and Strodthoff, 2023) backbone, and Diffusion-TS (Yuan and Qiao, 2024). For the proposed TimeBridge, we choose the data-dependent settings in Equation (19) and TimeBridge-GP denotes the time-dependent GP prior in Equation (20). As the prior mean and variance are derived from the training datasets, the proposed method does not require additional information.
Results are shown in Table 2. We observe that the proposed method outperforms the previous methods. The performance gain is significant in Context-FID and discriminative scores, reducing the scores of 40.70% and 18.05% on average, respectively. Notably, the strength of data and time-dependent prior is shown in real datasets, such as ETTh and Stocks. This indicates that our prior selection and modeling help in approximating the data distribution during synthesis. In average rank, the proposed method shows the best performance. It indicates the power of prior selection in time series generation within a similar structure.
We compare our model with Diffusion-TS on long-horizon generation tasks with a sequence length of 64. The results in Table 3 show that, on average, TimeBridge-GP produces higher-quality samples than standard Gaussian. Additional experiments with even longer sequences are reported in Appendix C.
@@ 6.3. Synthesis with Trend Condition
We now consider trend-guided generation, where the data samples are in $\mathcal{D}=\{({\mathbf{x}}^{i},{\mathbf{y}}^{i})\}_{i=1}^{N}$. For the trend ${\mathbf{y}}$, we consider three types: linear, polynomial with degree of three, and Butterworth filter (Butterworth et al., 1930), a signal processing filter that captures trends (Xia et al., 2024). Using these trends as conditions, we reimplement the experiments in Table 2 with the ETTh and Energy datasets. To apply the conditions to baseline methods, we feed the trend into a conditional embedding and concatenate it with the original embedding. For TimeBridge, we set the prior ${\mathbf{x}}_{T}={\mathbf{y}}$ upon the conditional embedding.
@@ 6.4. Imputation with Fixed Masks
For imputation, we compare the proposed method with imputation diffusion models and the conditional Langevin sampler of Diffusion-TS (Yuan and Qiao, 2024). We use random masks for the Mujoco dataset and the geometric mask from (Zerveas et al., 2021) for the ETTh and Energy datasets. We set the total length of imputation to 100 for the Mujoco dataset, matching the data length, and split the ETTh and Energy datasets into 48 time steps. For the proposed TimeBridge, we preserve the observed values and set the prior ${\mathbf{x}}_{T}$ as the interpolated values of the condition ${\mathbf{c}}$ as in Equation (21).
The imputation results are shown in Tables 5 and 6. Both tables demonstrate that TimeBridge achieves the lowest measures in imputation tasks. Since the Langevin sampler of Diffusion-TS does not require individual training for each missing ratio, we ensure a fair comparison by evaluating TimeBridge with the same training setting, TimeBridge-75%, which uses 75% missing values for training. TimeBridge-75% even outperforms the models trained with equal missing values, highlighting the generalization capability of TimeBridge in imputation tasks. The visualization of imputation results and sampling path can be found in Figures 5 and 4(b), respectively.
(a) Sampling of TimeBridge by applying the two pair-wise linear trends (red/blue) as priors.
@@
6.5. Additional Experiments
Ablation Study
Table 7 evaluates the effectiveness of training design with three variants: (1) variance exploding (VE) scheduler, (2) noise matching instead of directly predicting $D_{\theta}$ as in Zhou et al. (2024), and (3) without Fourier-based loss during training. The results indicate that each component of our diffusion scheduler and loss design contributes to improved performance.
Extended ablation results are provided in Appendix C.6.
Further Experiments
Please refer to Appendix C for further experimental results and findings, including the impact of prior selection measured by Wasserstein distance, evaluations on extreme data scenarios, and results of predictive score (Yoon et al., 2019).
6.6. Computation efficiency
Table 8 illustrates the strength of the proposed method, particularly when using the second-order Heun sampler during sampling. With fewer score function evaluations (NFE), our sampler achieves fast synthesis without compromising quality. We use 40 steps for sampling in all experiments, resulting in 119 NFEs with our settings. For comparison, we evaluate the Diffusion-TS using the same NFE. Given the trade-off of performance and sampling time in Diffusion-TS, our experimental results outperform previous methods. The marginal increase in time for TimeBridge is due to the time required to build the prior distribution.
@@ 6.6. Computation efficiency
Table 8. Comparison of sampling efficiency and generation performance. The time column indicates the generation time of each sample in milliseconds (ms) with a batch size of 512.
Data ETTh Stocks
Metric NFE Time C.-FID Disc. NFE Time C.-FID Disc.
Diffusion-TS 500 5.008 0.129 0.074 500 4.670 0.165 0.083
119 1.038 0.244 0.103 119 1.104 0.252 0.130
TimeBridge 119 1.426 0.067 0.052 119 1.429 0.079 0.052

### v3

@@ 4.1. Data- and Time-dependent Priors for Time Series
To investigate RQ1, we suggest using data prior to having a data-dependent distribution to better approximate the data distribution. Sampling from $\mathcal{N}(\mathbf{0},\mathbf{I})$ lacks data-specific information, which can be enhanced by enforcing that the priors capture data scale and temporal dependencies. Recently, the usages of data-dependent prior have been investigated for audio (popov2021grad; lee2022priorgrad) and image (yu2024constructing) domains. Therefore, we test the use of data-dependent prior for time series by setting the prior distribution as follows:
@@
In response to RQ2, we focus on capturing temporal dependencies in noise addition towards prior distribution. Recently, bilovs2023modeling injected stochastic processes instead of random noise to preserve the continuity of temporal data and han2022card utilized the pre-trained model for prediction and set the mean of prior as the output for regression.
As a solution, we employ Gaussian processes (GPs) for prior selection, which is effective for modeling temporal data (bilovs2023modeling; ansari2024chronos). GP is a distribution over function that defines the joint variability of an arbitrary pair of time points. For given sequence of timestamps $\Omega=\{1,\ldots,\tau\}$ and a function $\bm{f}_{GP}$ of timestamps that follows a GP with mean function $\bm{m}(\cdot)$ and the covariance function $\bm{\Sigma}(\cdot)$, we can draw ${\mathbf{x}}_{T}$ as follows:
@@ 4.2. Scale-preserving for Conditional Priors
To answer RQ3, we argue that the diffusion bridge is well-suited for preserving a given condition by setting the prior as the condition. As one of the common conditional setups, we consider the task of generating data that follows a given trend.
Conditional generation based on trend is straightforward by (i) setting the pair-wise trend to prior ${\mathbf{x}}_{T}={\mathbf{y}}$ and (ii) training the translation from trend to data. As the expected mean of ${\mathbf{x}}_{t}$ is a linear combination of ${\mathbf{x}}_{0}$ and ${\mathbf{x}}_{T}={\mathbf{y}}$ in Equation (8), the model learns the correction starting from the trend samples, utilizing the same data scale. In contrast, standard diffusion models are unsuitable for translating conditions into synthetic data; instead, they generate new data based on a specific trend condition, typically requiring additional steps such as providing guidance or correcting intermediate data samples ${\mathbf{x}}_{t}$ during sampling. Thus, we can also eliminate the need for penalty functions or corrections for sampling.
@@
The direct case of fixed point constraints is imputation. With the mask $\mathbf{m}$} indicating the missing value as 0, we construct the condition ${\mathbf{c}}={\mathbf{y}}\odot{\mathbf{m}}$ for imputation where $\odot$ indicates the Hadamard product. However, unlike soft constraints that apply to all values in ${\mathbf{y}}\in\mathbb{R}^{\tau\times d}$, the condition for imputation ${\mathbf{c}}\in\mathbb{R}^{\tau\times d}$ has missing values, which should be interpolated to data-dependent values rather than zeros. Therefore, we build local linear models for condition-to-time series imputation. We adopt linear spline interpolation (de1978practical) to generate a prior for fixed point constraints. In a feature-wise manner, let us consider $x\in\mathbb{R}$ as a single-dimensional input of ${\bm{x}}\in\mathbb{R}^{d}$. With a set $K$ of feature-wise observed times for $x$, the prior $x_{T}$ with the interpolation of each feature with timestamp $k$ is:
@@ 5. Related Works
Diffusion models outperform in time series synthesis tasks, compared to models based on VAE (naiman2024generative), GANs (yoon2019time). rasul2021autoregressive initially used diffusion networks for time series based on recurrent networks. For imputation, SSSD (lopez2023diffusionbased) and CSDI (tashiro2021csdi) considered time series imputation similar to image inpainting tasks. TimeDiff (shen2023non), LDT (feng2024latent), and TMDM (li2024transformermodulated) focused on time series forecasting tasks. coletta2024constrained investigated guiding the sampling with a penalty function called DiffTime. Our network design follows Diffusion-TS (yuan2024diffusionts), which introduced a decomposition architecture to disentangle time series data into trend and seasonal components and applied a Fourier-based loss term. Recent studies also explore frequency-domain diffusion (crabbe2024time) and transformer-based alternatives (chen2024sdformer).
@@
Several studies have brought Schrödinger bridges to time series tasks. chen2023provably analyzed algorithmic convergence and applied the bridge to imputation. hamdouche2023generative combined kernel-based bridges with feed-forward and LSTM networks for deep hedging. garg2024soft relaxed the Schrödinger bridge by setting a geometric mixture for the prior distribution and applied their method to time series. All of these methods require iterative solvers, which are computationally expensive and incompatible with standard diffusion pipelines. Also, none of them provides a unified framework for both unconditional and conditional generation. To the best of our knowledge, our paper is the first to use diffusion bridge, extending beyond classical Schrödinger bridge formulations and providing flexible priors.
@@
Table 2. Quality comparison of unconditional time series synthetic data with length of 24. The best results are in bold and the second best are underlined. We report the average score (lower is better) and the rank of each method. The baseline experimental results (denoted by $\dagger$) are adopted from (yuan2024diffusionts). We re-implement Diffusion-TS with their official code.
@@ 6.1. Experimental Setup
We assess the performance using widely used time series datasets. We use two simulation datasets: Sines with 5 features of different frequencies and phases of sine functions and MuJoCo of a multivariate advanced physics simulation with 14 features. For real-world datasets, we use ETT (Electricity Transformer Temperature) for long-term electric power with 7 features, Stocks of Google stock prices and volumes with 6 features, Energy, UCI dataset with appliances energy use in a low energy building with 28 features, and fMRI of the blood oxygen level-dependent functional MR imaging with 50 selected features.
For measures, we assess three metrics for unconditional time series generation on a normalized (0-1) scale. Context-Fréchet Inception Distance (Context-FID, C.-FID) score (jeha2022psa) measures the distributional quality for mean and variance projected on embedding on trained TS2Vec (yue2022ts2vec). Correlational (Corr.) score (liao2020conditional) measures temporal dependency of real and synthetic data using the absolute error of cross-correlation. Discriminative (Disc.) score (yoon2019time) measures the performance of a classifier trained to distinguish original and synthetic data in a supervised manner. For imputation, we use the mean squared error (MSE) and mean absolute error (MAE).
Our experiments are conducted using PyTorch on four NVIDIA GeForce RTX 4090 GPUs. We report the mean and its interval of five runs. For the details, refer to Appendix B.
@@
Table 5. Imputation results of MSE in the order of 1e-3 on the Mujoco dataset with a random mask of ratio {0.7, 0.8, 0.9}. The best results are in bold and the second best are underlined. The baseline experimental results are adopted from (lopez2023diffusionbased) and (yuan2024diffusionts).
@@ 6.2. Unconditional Generation
We demonstrate the unconditional generation tasks with a length of 24 as baselines. We compare with TimeVAE (desai2021timevae), TimeGAN (yoon2019time), Cot-GAN (xu2020cot), Diffwave (kong2021diffwave), DiffTime (coletta2024constrained) with SSSD (lopez2023diffusionbased) backbone, and Diffusion-TS (yuan2024diffusionts). For the proposed TimeBridge, we choose the data-dependent settings in Equation (18) and TimeBridge-GP denotes the time-dependent GP prior in Equation (19). As the prior mean and variance are derived from the training datasets, the proposed method does not require additional information.
Results are shown in Table 2. We observe that the proposed method outperforms the previous methods. The performance gain is significant in Context-FID and discriminative scores, reducing the scores of 40.70% and 18.05% on average, respectively. Notably, the strength of data and time-dependent prior is shown in real datasets, such as ETTh and Stocks. This indicates that our prior selection and modeling help in approximating the data distribution during synthesis. In average rank, the proposed method shows the best performance. It indicates the power of prior selection in time series generation within a similar structure. The results in Table 3 show that TimeBridge-GP produces higher-quality samples than standard Gaussian for long-horizon generation tasks with a sequence length of 64.
@@ 6.3. Synthesis with Trend Condition
We now consider trend-guided generation, where the data samples are in $\mathcal{D}=\{({\mathbf{x}}^{i},{\mathbf{y}}^{i})\}_{i=1}^{N}$. For the trend $\mathbf{y}$}, we consider three types: linear, polynomial with degree of three, and Butterworth filter (butterworth1930theory), a signal processing filter that captures trends (xia2024market). Using these trends as conditions, we reimplement the experiments in Table 2 with the ETTh and Energy datasets. To apply the conditions to baseline methods, we feed the trend into a conditional embedding and concatenate it with the original embedding. For TimeBridge, we set the prior ${\mathbf{x}}_{T}={\mathbf{y}}$ upon the conditional embedding.
@@ 6.4. Imputation with Fixed Masks
For imputation, we compare the proposed method with imputation diffusion models and the conditional Langevin sampler of Diffusion-TS (yuan2024diffusionts). We use random masks for the Mujoco dataset and the geometric mask from (zerveas2021transformer) for the ETTh and Energy datasets. We set the total length of imputation to 100 for the Mujoco dataset, matching the data length, and split the ETTh and Energy datasets into 48 time steps. For the proposed TimeBridge, we preserve the observed values and set the prior ${\mathbf{x}}_{T}$ as the interpolated values of the condition $\mathbf{c}$} as in Equation (20).
The imputation results are shown in Tables 5 and 6. Both tables demonstrate that TimeBridge achieves the lowest measures in imputation tasks. Since the Langevin sampler of Diffusion-TS does not require individual training for each missing ratio, we ensure a fair comparison by evaluating TimeBridge with the same training setting, TimeBridge-75%, which uses 75% missing values for training. TimeBridge-75% even outperforms the models trained with equal missing values, highlighting the generalization capability of TimeBridge in imputation tasks. The visualization of imputation results and sampling path can be found in Figures 6 and 5(b), respectively.
(a) Sampling of TimeBridge by applying the two pair-wise trends (red/blue) as priors.
@@
6.5. Additional Experiments
Ablation Study
Table 7 evaluates the effectiveness of training design with three variants: (1) variance exploding (VE) scheduler, (2) noise matching instead of directly predicting $D_{\theta}$ as in zhou2024denoising, and (3) without Fourier-based loss during training. The results indicate that each component of our diffusion scheduler and loss design contributes to improved performance.
Wasserstein distance
To experimentally justify the strength of the diffusion prior, we measure the Wasserstein distance within the diffusion bridge’s optimal transport framework. kollovieh2024flow argue that closer alignment of data and prior distributions reduces path complexity, and time-reflective priors can lower the Wasserstein distance. As shown in Table 8, our data-dependent priors (unconditional and imputation) yield shorter distances. While not always a direct indicator of performance, this may explain why diffusion bridge models surpass standard diffusion approaches.
Unconditional Imputation
ETTh 14.285 2.387 8.000 0.485
Energy 28.128 6.376 16.113 1.494
Extreme cases
Table 9. Ablation results on extreme values of top $k\%$ variance.
Top $k\%$ Metric ETTh Energy Stock
Diffusion-TS Ours Diffusion-TS Ours Diffusion-TS Ours
20% C.-FID 0.153 0.103 0.128 0.093 0.788 0.257
Corr. 0.095 0.070 0.946 0.150 1.217 0.010
50% C.-FID 0.144 0.098 0.139 0.096 0.475 0.154
Corr. 0.062 0.048 0.886 0.021 1.028 0.012
Longer Sequences
Modeling longer sequences is crucial for time series generation. Thus, we have tested lengths of 192 and 360, common for forecasting cyclic or annual patterns. By controlling only $\sigma_{data}$ and fixing other parameters to minimize the search space, Table 10 shows that our method outperforms standard diffusion priors, excelling in 11 out of 12 metrics.
Table 10. Comparison of longer sequences of 192 and 360.
Context-FID Cross-Correlation Discriminative
Stocks (192) 0.679 0.548 0.009 0.006 0.144 0.060
Stocks (360) 0.954 0.606 0.016 0.004 0.125 0.069
ETTh (192) 1.021 1.535 0.106 0.098 0.187 0.096
ETTh (360) 1.977 1.780 0.114 0.099 0.226 0.109
Quadratic Imputation Prior
[…]
6.6. Computation efficiency
Table 12 illustrates the computational time for constructing priors and training. For per-iteration overhead, one iteration of our method requires only $1.06\times$ the computation of Diffusion-TS. Moreover, during sampling, our sampler uses fewer NFEs and runs significantly faster.
Prior construction involves extracting only the empirical mean and variance from 18K training samples, which is comparable to model initialization time. Imputation priors require slightly more computation but remain marginal relative to full training cost. Linear spline priors are further accelerated using the Numba [^1] https://github.com/numba/numba package with JIT compilation.
Table 11. Comparison of linear and quadratic interpolation priors (MSE at different missing ratios).
Dataset Linear Quadratic
25% 50% 75% 25% 50% 75%
ETTh 0.66 0.20 0.15 0.63 0.20 0.12
Energy 12.97 23.07 63.84 15.62 30.05 62.42
Table 12. Detailed computation comparison for prior construction, training, and sampling.
Prior Construction Training Sampling
Method Init Uncond. Imput. Iter/step Total Time
Diffusion-TS 0.18s 0.081s 1458s 90.14s
[…]